# Beyond Linear Probes: Dynamic Safety Monitoring for Language Models

### v1

@@ 1 Introduction
However, despite widespread benefits to model capabilities, we have yet to see dynamic computation (Han et al., 2021) bring similar flexibility to AI safety. This is particularly true in the domain of LLM safety monitors, trained to detect harmful requests (Han et al., 2024), or problematic model behavior (Goldowsky-Dill et al., 2025; MacDiarmid et al., 2024; Chaudhary & Barez, 2025). Popular monitoring techniques include LLM-as-judges of natural language on the one hand (Inan et al., 2023; Zeng et al., 2025), and cheap linear probes in activation-space on the other (Alain & Bengio, 2017). In both cases, we argue that current approaches are inflexible. Considering that most requests are benign, dedicated LLMs have an excessively large minimum cost as always-on monitors, while activation probes provide only the most basic, static guardrails. In contrast, activation monitors that scale dynamically offer two key benefits:
@@ 3.2.2 Cascading defense
Concretely, one first evaluates the linear probe $y=P^{[N]}_{:1}(\bm{z})=w^{[0]}+{\bm{z}^{\top}\bm{w}^{[1]}}\in\mathbb{R}$. Additional higher-order terms are then evaluated only if the prediction falls outside a confidence threshold $\sigma(y)\notin(\tau,1-\tau)$. This cascading polynomial defense is described in Algorithm 1.
@@ Acknowledgments
We are grateful to Jakub Vrábel, Tung-Yu Wu, Roy Miles, Grigorios Chrysos, Mihalis Nicolaou, Zhi-Yi Chin, Jaeyoung Lee, Adi Simhi, and Vincent Wang for feedback on earlier drafts and/or helpful discussions throughout the project.
@@ Appendix B PyTorch implementation
Here we provide a simple PyTorch implementation of the truncated polynomials in Eq. 5. Assuming the relevant weight matrices are defined upon initialization, TPCs’ forward pass can be implemented straightforwardly via Fig. 6.
Figure 6: Truncated polynomial forward pass
def forward(self, x, test_time_order):
# linear probe
y = einsum(self.W[1], x, o i, i -> o’) + self.W[0]
[…]
for n in range(min(test_time_order, self.max_order)-1):
order = n+2
inner = einsum(x, self.HO[n], ’… i, r i -> r’) ** (order)
yn = einsum(inner, self.lam[n], ’… r, r -> …’)
[…]
@@ G.1 Progressive training ablations
We first perform additional ablations on the proposed progressive training scheme, across all $4$ models, on both datasets. We show the F1 scores evaluating truncated models with and without the proposed progressive training in Figs. 11 and 12. As can be seen, the proposed progressive scheme leads to truncated models performing much better than with regular training.
@@ G.3 Rank ablations
We also fully tune and re-train all baseline models for the various other choices of rank $R$ for additional comparisons and ablations alike, which are shown in Figs. 18, 19 and 20.

### v2

@@ 1 Introduction
However, despite widespread benefits to model capabilities, dynamic computation (Han et al., 2021) for AI safety remains nascent. This is particularly true in the domain of LLM safety monitors, trained to detect harmful requests (Han et al., 2024), or problematic model behavior (Goldowsky-Dill et al., 2025; MacDiarmid et al., 2024; Chaudhary & Barez, 2025). Popular monitoring techniques include LLM-as-judges of natural language on the one hand (Inan et al., 2023; Zeng et al., 2025), and cheap linear probes in activation-space on the other (Alain & Bengio, 2017). In both cases, we argue that current approaches are inflexible. Considering that most requests are benign, dedicated LLMs have an excessively large minimum cost as always-on monitors, while activation probes provide only the most basic, static guardrails. Whilst recent work proposes to chain these two existing approaches (McKenzie et al., 2025; Cunningham et al., 2025), requiring large external LLMs that need finetuning/prompting limits their flexibility. In contrast, a single activation monitor that scales dynamically offers two key benefits:
@@ 2 Related work
Cascades & ensembles
Combining or learning multiple submodels is a powerful way to improve upon single models. Multiple classifiers used in a cascade (or networks with early exits) bring robustness and/or speed in computer vision (Viola & Jones, 2004; Bourdev & Brandt, 2005; Romdhani et al., 2001), machine learning (Grubb & Bagnell, 2012; Xu et al., 2012), and deep neural networks (Teerapittayanon et al., 2016; Raposo et al., 2024; Yue et al., 2024) alike. Recent work similarly explores the combination of multiple models for LLM monitoring (McKenzie et al., 2025; Hua et al., 2025; Cunningham et al., 2025). Concretely, McKenzie et al. (2025); Cunningham et al. (2025) both show large computational savings using activation probes as a first line of two-stage defense–routing inputs to external LLM-as-monitors when uncertain. Whilst well-positioned to benefit from future LLM advances, both McKenzie et al. (2025); Cunningham et al. (2025) require additional LLM fine-tuning or prompting, and calls to extra LLMs during inference time. Instead, TPCs learn dynamic $N$-layer defense from neuron interactions, directly in the original LLMs’ activation space–offering built-in neuron attribution. We view these methods as complementary; in principle, a cascade of depth $N+1$ could combine TPCs with an LLM-as-monitor final layer for additional defense.
@@ 3.2.2 Cascading defense
Similar to recent work (McKenzie et al., 2025; Cunningham et al., 2025), we first evaluate the linear probe $y=P^{[N]}_{:1}(\bm{z})=w^{[0]}+{\bm{z}^{\top}\bm{w}^{[1]}}\in\mathbb{R}$. We then add additional higher-order terms only if the prediction falls outside a confidence threshold $\sigma(y)\notin(\tau,1-\tau)$. This cascading polynomial defense is described in Algorithm 1.
@@ Acknowledgments
We are grateful to Jakub Vrábel, Tung-Yu Wu, Roy Miles, Grigorios Chrysos, Mihalis Nicolaou, Zhi-Yi Chin, Jaeyoung Lee, Adi Simhi, and Vincent Wang for feedback on earlier drafts and/or helpful discussions throughout the project. We also thank Dmitrii Krasheninnikov for helpful pointers to recent related work.
@@ Appendix B PyTorch implementation
Here we provide a simple PyTorch implementation of the truncated polynomials in Eq. 5. Assuming the relevant weight matrices are defined upon initialization, TPCs’ forward pass can be implemented straightforwardly via LABEL:lst:poly-forward.
Listing 1: Truncated polynomial forward pass
1def forward(self, x, test_time_order):
2 # linear probe
3 y = einsum(self.W[1], x, 'o i, ... i -> ... o') + self.W[0]
4
5 # loop over higher-orders
6 for n in range(min(test_time_order, self.max_order)-1):
7 order = n+2
8 inner = einsum(x, self.HO[n], '... i, r i -> ... r') ** (order)
9 yn = einsum(inner, self.lam[n], '... r, r -> ...')
10
11 y = y + yn # add nth component
@@ G.1 Progressive training ablations
We first perform additional ablations on the proposed progressive training scheme, across all $4$ models, on both datasets. We show the F1 scores evaluating truncated models with and without the proposed progressive training in Figs. 10 and 11. As can be seen, the proposed progressive scheme leads to truncated models performing much better than with regular training.
@@ G.3 Rank ablations
We also fully tune and re-train all baseline models for the various other choices of rank $R$ for additional comparisons and ablations alike, which are shown in Figs. 17, 18 and 19.